OpenStudy (anonymous):

find the inverse of the function F(x)= 1/6x+7

OpenStudy (jdoe0001):

\(\bf f(x)=\color{red}{y}= \cfrac{1}{6x}+7\qquad inverse\implies x=\cfrac{1}{6\color{red}{y}}+7\) notice to find the inverse function, all you do is swap the variables about then solve for "y"

OpenStudy (anonymous):

all you do is switch x and y and since f(x) is basically y you would have x=1/6y+7. then you have to bring the 7 over to the other side and you would get x-7=1/6y. divide by 1/to get y by itself so the answer would be y=6x-42

OpenStudy (anonymous):

divide by 1/6 *

OpenStudy (anonymous):

i got it :)

OpenStudy (anonymous):

thank you :)